OpenStudy (anonymous):

A 15.0 cm long cylindrical glass tube, sealed at one end, is filled with ethanol. The mass of ethanol needed to fill the tube is found to be 11.86 g. The density of ethanol is 0.789 g/mL. Calculate the inner diameter of the tube in cm.

OpenStudy (anonymous):

Please help me with my another problem

OpenStudy (anonymous):

Thank you again

OpenStudy (anonymous):

I have lots of problem here qsx. Can you help me? please

OpenStudy (anonymous):

when is all due? i can help you with one more i have pre-cal exam tomorrow

OpenStudy (anonymous):

The US quarter has a mass of 5.67 g and his approximately 1.55mm thick. (A) how many quarters would have to be stacked to reach 575 ft. what is the height of the washington monument? (B) how much would this weigh?(C) how much money would this stack contain(D) at the begging of 2007, the natl. depth was $8.7 trillion. How many stack like one describe would be necessary to pay off this debt?

OpenStudy (anonymous):

That's the problem qsx. That's the last one.

OpenStudy (anonymous):

lol piece of cake, wait 5 mins

OpenStudy (anonymous):

Ok, i can wait

OpenStudy (anonymous):

Hey, i need to go. I'm already late for my class, thank you for helping. ^_^ bye

OpenStudy (anonymous):

np prob, i am still solving

OpenStudy (anonymous):

@Festinger please help me with this problem

OpenStudy (festinger):

1 foot = 304.8mm \[575feet * \frac{304.8}{1 foot}= 175260mm\] this would be: \[\frac{175260mm}{1.55mm/coin}= 113071 coins\] and thus \[113071*5.67=641112g=641kg\]

OpenStudy (anonymous):

The US quarter has a mass of 5.67 g and his approximately 1.55mm thick. (A) how many quarters would have to be stacked to reach 575 ft. what is the height of the washington monument? (B) how much would this weigh?(C) how much money would this stack contain(D) at the begging of 2007, the natl. depth was $8.7 trillion. How many stack like one describe would be necessary to pay off this debt? @Festinger i mean with this problem. I'm so sorry :33

OpenStudy (anonymous):

I gave you the wrong one. I'm so sorry T_T

OpenStudy (festinger):

(c) it's worth whatever 113071 quarters is worth in dollars. (d) take $8.7*10^12 divided by whatever one quarter is worth in $ and multiply by 1.55mm

OpenStudy (anonymous):

Ohh. Thank you so much ^_^

OpenStudy (anonymous):

By the way, how to solve the problem in letter B?

OpenStudy (festinger):

Take the number of coins times the weight of each coin. It's in one of my replies. I solved (a) and (b) in one reply, the (C) and (D) in the second.

OpenStudy (anonymous):

Okaay. :) thank you so much festinger. ^_^ you really helped me a lot.